Write the equation of the line that is parallel to the Line 4y+3x-21=0 and passes through the point (16, -5)
Answer by stanbon(75887)

Write the equation of the line that is parallel to the Line 4y+3x-21=0 and passes through the point (16, -5)
----
Find the slope of the given line::
4y+3x-21 = 0
4y = -3x+21
y = (-3/4)x + (21/4)
slope = -3/4
------------------------
Form of your equation:: y = mx + b
---
Solve for "b" when x = 16, y = -5, and m = -3/4
------
-5 = (-3/4)16 + b
-5 = -12 + b
b = 7
------
Equation:
y = (-3/4)x + 7
--------------------
Cheers,
Stan H.
-------------